OpenStudy (anonymous):

Given the explicit formula for an arithmetic sequence, find the first five terms and the 10th term. an = -11 + 7n

OpenStudy (anonymous):



OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

well, you have everything, just set n = 1... 5 and 10

OpenStudy (anonymous):

@Skaternerd123 is a bad guy

OpenStudy (jdoe0001):

\(\bf \Large a_n = a_1+(n-1)d\\ a_n = -11 + 7n\\ \textit{to find term "whatever"}\\ \Large a_{whatever} = -11 + 7(whatever)\)

OpenStudy (anonymous):

@juliorrethatoku please do not tag my name smelly boy

OpenStudy (anonymous):

Lol

OpenStudy (anonymous):

please leave me alone bully I am reporting you!! @Skaternerd123